# Dissecting the NVIDIA Hopper Architecture through Microbenchmarking and Multiple Level Analysis

### v1

@@ 6.2 Instruction-level
TABLE XIII: Power consumption and energy efficiency of maximum shape under wgmma instructions. Frequency is measured in MHz. Performance is measured in TFLOPS (TOPS). Efficiency is measured in TFLOPS (TOPS)/watt.
@@
6.3 Library-level
The Transformer Engine (TE) [43] is a library specifically designed to accelerate Transformer models [44], following the introduction of the Hopper architecture. It is capable of leveraging the FP8 precision offered by both the Hopper and Ada architectures. In particular, it provides a variety of optimized modules for Transformer layers that can be utilized within the widely-used deep learning framework, PyTorch [45]. In this subsection, we benchmark two modules of the Transformer Engine to explore their potential as key components in large language models.
6.3.1 Linear Layer
[…]
As depicted in Fig. 6, when performing matrix multiplication with relatively small matrix sizes, the proportion of overhead attributed to conversion is significantly larger than that resulting from the GEMM kernel computation in FP8 Tensor Cores. To evaluate and investigate the support and optimization of TE for linear layers, we measure the throughput (GFLOPS) of te.Linear for square matrices $\mathbf{D}_{n\times n}=\mathbf{A}_{n\times n}\times\mathbf{B}_{n\times n}$. We choose square matrices for their simplicity while still effectively demonstrating the characteristics of te.Linear.
We extensively assess the Linear performance across diverse shapes, data types, and hardware setups (Fig. 7). Leveraging the Transformer Engine, we expedite matrix multiplications for the Linear layer using FP8 Tensor Cores. Our findings reveal an increase in GPU utilization and throughput with larger matrix sizes. FP8 performance is influenced by the overhead from data format conversion and quantization operators. For smaller matrix sizes, FP8 throughput is lower compared to FP16 or FP32. However, with $N$=8192, FP8’s performance gains become evident. When $N$=16384, H800 and 4090 utilizing FP8 achieve almost twice the throughput of FP16. This underscores FP8’s high throughput potential but underlines the need for specific conditions to attain optimal computing density.
Figure 6: Proportion of execution time for different operators when performing FP8 matrix multiplication using te.Linear.
6.3.2 Transformer Layer
@@
TE offers a te.TransformerLayer module that encompasses all operator optimizations for transformer layer structures, facilitating the implementation of various Large Language Model (LLM) structures by adjusting its parameters. However, some operators, such as Softmax and GeLU, have not been quantized to FP8 by TE, resulting in significant data format conversion overhead. Additionally, the DotProductAttention operator, implemented with flash-attention [46], does not utilize FP8 Tensor Cores.
The computational overhead of the transformer layer’s linear layer primarily depends on the hidden size, raising the question of which hidden state (dimension of embedding) will yield better performance for TE with FP8 compared to FP16. We investigate this by examining the open-source LLM, Llama [47, 48], modifying the activation function to SwiGLU [49] and normalization to RMSNorm [50]. We set layer structure parameters based on the hidden state’s size, with hidden states 4096, 5120, and 8192 corresponding to Llama configurations 7b, 13b, and 70b, respectively.
TABLE XIV: Parameter settings of te.TransformerLayer for various $hidden\_size$ values
@@
We fixed the input as (4, 512, $hidden\_size$), where 4 is the $batch\_size$, 512 is the $sequence\_length$, and the attention mask is set to None. We then calculated the latency (ms) required for encoding a single layer once, focusing on the encoding task for a single layer.
The Transformer Engine condenses the entire Transformer Layer structure into te.TransformerLayer. Fig. 8 illustrates the latency for the same input text, offering a performance comparison across various hardware setups and data types with te.TransformerLayer. As computational density increases, the advantage of H800 in computation becomes evident. Notably, FP16 shows nearly twice the speed compared to FP32. FP8 outperforms FP16 for hidden_size$>$4096 but does not achieve double FP16 performance. This is because some modules within the Transformer Layer still do not utilize FP8 precision for calculations and data movement.
Figure 7: Comparison of throughput for matrix multiplication with two same-size matrices $\mathbf{D}_{n\times n}=\mathbf{A}_{n\times n}\times\mathbf{B}_{n\times n}$ in different hardware configurations and data types using te.Linear.
Figure 8: Comparison of latency for the same input text in different hardware configurations and data types using te.TransformerLayer.
6.4 Application-level
Currently, the Transformer Engine has not provided optimal support for mainstream decode-only casual language models. In order to test the inference performance of the Transformer Engine on this type of model like, Llama [47], we replaced the nn.Linear and RMSNorm in the original model structure with te.Linear and te.RMSNorm, respectively, to ensure that most modules in the model utilize the Transformer Engine.
To evaluate the effectiveness of TE in generating text for Llama, we follow [51] using the ShareGPT dataset as input for the LLM. The ShareGPT dataset comprises conversations between users and ChatGPT [52], which have been shared by the users. We tokenize these datasets and, based on their input and output lengths, generate synthesized client requests.
@@ 6.4 Application-level
We use throughput as the evaluation metric, which represents the total text length that can be processed per second: $Throughput=(input\_len+output\_len)/time$.
LLM inference throughput results. We test the state-of-the-art decode-only models on inference with different data types, as shown in Table XV. We set the input length and text generation length to be relatively short, and the decode-only model is memory-bound during inference, so the computational advantages of FP8 Tensor Cores are not significant. Moreover, since the current Transformer Engine does not provide comprehensive support, data transmission between modules still occurs in FP16/FP32 without operator fusion. It is possible that when the model size and input data length increase, and with good operator fusion support, a certain improvement can be achieved.
TABLE XV: Inference Throughput (Tokens / s) for different model sizes on different GPUs and different data types
@@
7 Distributed Shared Memory
@@ 7.1 Latency
As shown in Table III, when we directly access the shared memory, the latency is 29 cycles. However, when we access the block’s local shared memory by the distributed shared memory interface, the latency reaches 33 cycles. The results indicate that even if we access the local shared memory and it is not required to leverage SM-to-SM network, the distributed shared memory still introduces some overhead. When the cluster size increases to two, the access latency between SMs is 181 cycles, a 32% reduction compared to L2 cache. Notice that transferring data from one SM to another via global memory incurs a cost of 956 cycles (one store and one load), utilizing DSM can reduce this latency by 5.28$\times$, which is close to the 7$\times$ reduction reported in NVIDIA’s official documentation. When the cluster size increases from two to sixteen (the maximum cluster size in Hopper), the access latencies between different blocks range from 184 to 213 cycles.
7.2 Throughput
@@
[…]
Figure 9: Scenarios for DSM throughput testing
For the inter-SM case, since DSM allows any block within a cluster to access another block, the access pattern significantly affects data movement throughput. Thus, as shown in Fig. 9, we designed three different copy scenarios to evaluate DSM throughput. The first is a ring-based access scheme, which is common in high-performance computing (e.g., ring allreduce [53, 54], matrix multiplication [55, 56]). In each cluster, threads in a block ranked by $R$ access the shared memory of the block ranked by $(R+1)\%CS$, where $CS$ is the cluster size. The second is pair-based, often used in butterfly communication scenarios in high-performance computing. The third is broadcast-based, a common access pattern in various parallel algorithms. In additon to above access pattern, we are also trying to explore the block scheduling policy within the cluster. There are three scheduling policies available in Hopper: default, Spread, and LoadBalancing. We explore the throughput of different access patterns under each scheduling policy. In the remaining experiments in this section, we launch 50160 blocks to ensure sufficient block residency on the SMs. Additionally, we allocate 16KB of shared memory for each block.
@@
Figure 10: Performance of access pattern on different policies
@@
Figure 11: The data communication throughput of the SM-to-SM network. “ILP” refers to the number of parallelizable data movement instructions.
7.3 Application-level
@@
We select histogram as an application-level benchmark. Histograms are frequently employed in image processing and data mining to analyze the distribution of data elements by displaying the occurrence frequency of each element’s value. While GPUs can accelerate histogram computation, achieving high computational efficiency is challenging. We redesign the histogram [^3] https://github.com/NVIDIA/cuda-samples/Samples/2_Concepts_and_Techniques/histogram application using DSM, distributing bins across blocks in the same cluster. During histogram counting via shared memory, each thread loads the element and determines the DSM address for the target bin, followed by an atomic increment operation. We adjust cluster size, block size, and bin count, measuring element processing throughput.
@@
Figure 12: Performance of the histogram application with distributed shared memory. The throughput is measured by the number of processing elements per second. “CS” refers to the cluster size. “Nbins” refers to the number of histogram bins.
Fig. 12 displays the histogram throughput with distributed shared memory. First, the optimal cluster size differs for various block sizes (CS=4 for block size 128, CS=2 for block size 512). Increasing block and cluster sizes can saturate SM-to-SM network utilization, potentially degrading overall performance due to resource contention. Second, a notable performance drop occurs from 1024 to 2048 Nbins when CS=1 (equivalent to not utilizing DSM). Larger Nbins demand more shared memory space and limit active block numbers on an SM. Employing the cluster mechanism to divide Nbins within the same cluster enhances block concurrency, mitigating this issue. Lastly, although shared memory is not a limiting factor for active block numbers with block size = 512, choosing an appropriate cluster size eases the on-chip shared memory traffic by leveraging the SM-to-SM network resource, ultimately improving overall performance.
8 Dynamic Programming Instructions
@@
Nvidia offers DPX functions [^4] https://docs.nvidia.com/cuda/cuda-c-programming-guide/index.html#dpx from CUDA 12 onward to accelerate dynamic programming code, enhancing programming ease. On the latest Hopper architecture, these functions are hardware-accelerated. In this section, we conduct both instruction-level and application-level benchmarks to comprehensively evaluate the potential and real-world performance of DPX.
8.1 Instruction-level
@@
Fig. 13 and Fig. 14 show the latency and throughput of the DPX functions on three tested GPUs. Since the DPX of RTX4090 and A100 are software emulation, their performance is almost the same. What can be observed is that for relu instructions, the performance of H800 is significantly better than the other two. For 16-bit operations, H800 also has significant acceleration, up to 13 times.
@@
Figure 13: Latency of DPX functions on different devices
Figure 14: Throughput of DPX functions on different devices
8.2 Application-level
This case study explores the application of DPX instructions to the Smith-Waterman algorithm[57], a classic dynamic programming algorithm used in bioinformatics. The Smith-Waterman algorithm finds optimal local alignments between two sequences and has wide-ranging applications in sequence alignment, genome analysis, and protein structure prediction. It employs dynamic programming by filling a two-dimensional matrix to compute the optimal local alignment score. Each matrix element represents the optimal local alignment score at corresponding positions within the two sequences. The matrix is filled recursively using a specific formula, ultimately yielding the optimal alignment.
@@ 8.2 Application-level
Our testing methodology is based on [58]. We use the same code, data, and testing parameter configuration as described in [58]. We use 20 protein sequences, with lengths ranging from 144 to 5478, as queries against a simulated databases. The database contains one million sequences, with fixed lengths 1024. Our results represent the average performance across all twenty queries. The performance metric for our experiments is GCUPS (Giga Cell Updates Per Second), representing the number of dynamic programming matrix cell updates performed per second.
Figure 15: Performance of the Smith-Waterman algorithm across various data types
@@
9 Conclusion
This paper examines the memory hierarchy and tensor core performance of the latest three Nvidia GPU architectures using instruction-level benchmarks. We found that the Hopper architecture demonstrates advantages in memory bandwidth and tensor core performance, aligning with official claims. Additionally, the impact of the L2 partitioned cache on program performance in Ampere and Hopper should be noted. Utilizing the latest wgmma instructions is essential to fully leverage the fourth-generation tensor core’s capabilities. We analyze AI performance across various architectures at both library and application levels, highlighting the impact of different precisions. Experiments indicate that low-precision data types offer greater advantages with larger operation scales. Furthermore, we explore key features of the Hopper architecture. Distributed shared memory exhibits varying performance across different access patterns and scheduling policies, necessitating careful parameter consideration in program design. TMA does not affect memory bandwidth but requires attention to operation shape. We discovered that DPX does not always enhance performance. Our research improves understanding of the latest architecture’s characteristics and performance, aiding in optimized algorithm design and application performance.
Acknowledgments

### v2

@@ 6.2. Instruction-level
Table 11. Power consumption and energy efficiency of maximum shape under wgmma instructions. Frequency is measured in MHz. Performance is measured in TFLOPS (TOPS). Efficiency is measured in TFLOPS (TOPS)/watt.
@@
6.3. Library-level
The Transformer Engine (TE) (NVIDIA, 2022) is a library specifically designed to accelerate Transformer models (Vaswani et al., 2017), following the introduction of the Hopper architecture. It is capable of leveraging the FP8 precision offered by both the Hopper and Ada architectures. In particular, it provides a variety of optimized modules for Transformer layers that can be utilized within the widely-used deep learning framework, PyTorch (Paszke et al., 2019). In this subsection, we benchmark two modules of the Transformer Engine to explore their potential as key components in large language models.
6.3.1. Linear Layer
In the Transformer architecture, most of the computational overhead comes from the linear layers, specifically matrix multiplication while the Transformer Engine provides the te.Linear implementation to perform matrix multiplication with higher throughput on FP8 Tensor Cores. When employing the Transformer Engine with te.Linear for matrix multiplication in FP8 precision, TE converts both the input and weights in the linear layer to FP8. This conversion process involves data transformation and quantization operations. For example, as the dynamic range of FP8 may not encompass the maximum value of the input tensor, TE identifies the maximum absolute value of the input data as the scaling factor. It then adjusts the input data to fit the representation range of FP8 using $\mathit{inp_{fp8}}=\mathit{inp_{fp16}}/\mathit{scale}$, followed by matrix multiplication in FP8 Tensor Core $\mathit{out_{fp8}}=\mathit{inp_{fp8}}\times\mathit{w_{fp8}}$. Finally, it scales the result with $\mathit{out_{fp16}}=\mathit{out_{fp8}}\times\mathit{scale}$. This operation would introduce some overhead.
When performing FP8 matrix multiplication using te.Linear, kernel execution time significantly increases with matrix size. For smaller matrices (N=4096), kernel time accounts for only 25.3% of the execution, while for larger matrices (N=16384), it dominates at 84.7%. To evaluate Tensor Engine (TE) optimization for linear layers, we measure the throughput (GFLOPS) of te.Linear using square matrices $\mathbf{D}_{n\times n}=\mathbf{A}_{n\times n}\times\mathbf{B}_{n\times n}$, effectively showcasing its performance characteristics.
Figure 6. Comparison of throughput for matrix multiplication with two same-size matrices $\mathbf{D}_{n\times n}=\mathbf{A}_{n\times n}\times\mathbf{B}_{n\times n}$ in different hardware configurations and data types using te.Linear.
We extensively assess the Linear performance across diverse shapes, data types, and hardware setups (Fig. 6). Leveraging the Transformer Engine, we expedite matrix multiplications for the Linear layer using FP8 Tensor Cores. Our findings reveal an increase in GPU utilization and throughput with larger matrix sizes. FP8 performance is influenced by the overhead from data format conversion and quantization operators. For smaller matrix sizes, FP8 throughput is lower compared to FP16 or FP32. However, with NN=8192, FP8’s performance gains become evident. When NN=16384, H800 and 4090 utilizing FP8 achieve almost twice the throughput of FP16. This underscores FP8’s high throughput potential but underlines the need for specific conditions to attain optimal computing density.
6.3.2. Transformer Layer
@@
TE offers a te.TransformerLayer module that encompasses all operator optimizations for transformer layer structures, facilitating the implementation of various Large Language Model (LLM) structures by adjusting its parameters. However, some operators, such as Softmax and GeLU, have not been quantized to FP8 by TE, resulting in significant data format conversion overhead. Additionally, the DotProductAttention operator, implemented with flash-attention (Dao et al., 2022), does not utilize FP8 Tensor Cores.
The computational overhead of the transformer layer’s linear layer primarily depends on the hidden size, raising the question of which hidden state (dimension of embedding) will yield better performance for TE with FP8 compared to FP16. We investigate this by examining the open-source LLM, Llama (Touvron et al., 2023a, b), modifying the activation function to SwiGLU (Shazeer, 2020) and normalization to RMSNorm (Zhang and Sennrich, 2019). We set layer structure parameters based on the hidden state’s size, with hidden states 4096, 5120, and 8192 corresponding to Llama configurations 7b, 13b, and 70b, respectively.
Table 12. Parameter settings of te.TransformerLayer for various hidden_size values
@@
We fixed the input as (4, 512, hidden_sizehidden\_size), where 4 is the batch_sizebatch\_size, 512 is the sequence_lengthsequence\_length, and the attention mask is set to None. We then calculated the latency (ms) required for encoding a single layer once, focusing on the encoding task for a single layer.
Figure 7. Comparison of latency for the same input text in different hardware configurations and data types using te.TransformerLayer.
The Transformer Engine condenses the entire Transformer Layer structure into te.TransformerLayer. Fig. 7 illustrates the latency for the same input text, offering a performance comparison across various hardware setups and data types with te.TransformerLayer. As computational density increases, the advantage of H800 in computation becomes evident. Notably, FP16 shows nearly twice the speed compared to FP32. FP8 outperforms FP16 for hidden_size>>4096 but does not achieve double FP16 performance. This is because some modules within the Transformer Layer still do not utilize FP8 precision for calculations and data movement.
6.4. Application-level
Currently, the Transformer Engine has not provided optimal support for mainstream decode-only casual language models. In order to test the inference performance of the Transformer Engine on this type of model like, Llama (Touvron et al., 2023a), we replaced the nn.Linear and RMSNorm in the original model structure with te.Linear and te.RMSNorm, respectively, to ensure that most modules in the model utilize the Transformer Engine.
To evaluate the effectiveness of TE in generating text for Llama, we follow (Kwon et al., 2023) using the ShareGPT dataset as input for the LLM. The ShareGPT dataset comprises conversations between users and ChatGPT (OpenAI, [n. d.]), which have been shared by the users. We tokenize these datasets and, based on their input and output lengths, generate synthesized client requests.
@@ 6.4. Application-level
We use throughput as the evaluation metric, which represents the total text length that can be processed per second: Throughput=(input_len+output_len)/timeThroughput=(input\_len+output\_len)/time.
Table 13. Inference Throughput (Tokens/s) for different model sizes on different GPUs and different data types
@@
LLM inference throughput results. We test the state-of-the-art decode-only models on inference with different data types, as shown in Table 13. We set the input length and text generation length to be relatively short, and the decode-only model is memory-bound during inference, so the computational advantages of FP8 Tensor Cores are not significant. Moreover, since the current Transformer Engine does not provide comprehensive support, data transmission between modules still occurs in FP16/FP32 without operator fusion. It is possible that when the model size and input data length increase, and with good operator fusion support, a certain improvement can be achieved.
Insight 1. wgmma can achieve 95% of Hopper’s theoretical peak performance, while the backward-compatible mma can only reach 62.9%. 2. Using wgmma requires attention to power wall constraints, whereas mma demonstrates better energy efficiency. 3. Utilizing 4th-generation Tensor Core low-precision formats (FP8) is beneficial for libraries and applications, but careful consideration must be given to problem size, precision conversion overhead and accuracy loss.
7. Distributed Shared Memory
@@ 7.1. Latency
[…]
7.2. Throughput
@@
For the intra-SM case, the theoretical throughput of shared memory per SM is calculated as $1755\,\text{MHz}\times 128\,\text{Bytes}=225\,\text{GB/s}$. Using the DSM interface to access the block’s own memory only reaches 205 GB/s, 80% of peak performance. It is noteworthy that intra-SM shared memory access can also be performed without using the DSM interface, achieving over 99.8% of theoretical performance, as illustrated in Table 5. This indicates that there exists an additional overhead of accessing the block’s shared memory through DSM.
Figure 8. Scenarios for DSM throughput testing
For the inter-SM case, since DSM allows any block within a cluster to access another block, the access pattern significantly affects data movement throughput. Thus, as shown in Fig. 8, we designed three different copy scenarios to evaluate DSM throughput. The first is a ring-based access scheme, which is common in high-performance computing (e.g., ring allreduce (Thakur et al., 2005; Sergeev and Del Balso, 2018), matrix multiplication (Cannon, 1969; Chtchelkanova et al., 1997)). In each cluster, threads in a block ranked by $R$ access the shared memory of the block ranked by $(R+1)\%CS$, where $CS$ is the cluster size. The second is pair-based, often used in butterfly communication scenarios in high-performance computing. The third is broadcast-based, a common access pattern in various parallel algorithms. In addition to above access pattern, we are also trying to explore the block scheduling policy within the cluster. There are three scheduling policies available in Hopper: default, Spread, and LoadBalancing. We explore the throughput of different access patterns under each scheduling policy. In the remaining experiments in this section, we launch 50160 blocks to ensure sufficient block residency on the SMs. Additionally, we allocate 16KB of shared memory for each block.
@@
Figure 9. Performance of access pattern on different policies
@@
Figure 10. The data communication throughput of the SM-to-SM network. “ILP” refers to the number of parallelizable data movement instructions.
7.3. Application-level
@@
We select histogram as an application-level benchmark. Histograms are frequently employed in image processing and data mining to analyze the distribution of data elements by displaying the occurrence frequency of each element’s value. While GPUs can accelerate histogram computation, achieving high computational efficiency is challenging. We redesign the histogramhttps://github.com/NVIDIA/cuda-samples/Samples/2_Concepts_and_Techniques/histogram application using DSM, distributing bins across blocks in the same cluster. During histogram counting via shared memory, each thread loads the element and determines the DSM address for the target bin, followed by an atomic increment operation. The histogram application has the following two characteristics: 1) The data access flow is many-to-one similar to broadcasting (shared memory in each block is accessed by multiple blocks). 2) Due to the randomness of input data, writes to shared memory may generate unpredictable bank conflicts, which keeps ILP at a relatively low level. We adjust cluster size, block size, and bin count, measuring element processing throughput.
@@
Figure 11. Performance of the histogram application with distributed shared memory. The throughput is measured by the number of processing elements per second. “CS” refers to the cluster size. “Nbins” refers to the number of histogram bins.
The results shown in Fig. 11 can be summarized in two points. First, in addition to cluster size from 1 to 2, as cluster size increases, performance actually decreases. This reflects the broadcast-like access pattern degrading performance with larger clusters, corresponding to the results in Fig. 9. Second, increasing block size can significantly improve throughput. This demonstrates how higher block sizes can compensate for low ILP scenarios, which corresponds to the results in Fig. 10. These findings demonstrate how the fundamental DSM behavior patterns we identified can be used to predict and understand real application performance, providing practical guidance for optimizing DSM-based applications.
Beyond the instruction-to-application analysis, we observe that the optimal cluster size differs for various block sizes (CS=4 for block size 128, CS=2 for block size 512). This suggests that kernel configuration parameters should be carefully selected to maximize performance. Additionally, when using DSM, we achieve over 30% performance improvement compared to not using it (CS=1 for block size 512). Overall, choosing an appropriate cluster size eases the on-chip shared memory traffic by leveraging the SM-to-SM network resource, ultimately improving overall performance.
Insight 1. DSM reduces inter-SM communication latency by 6.1×. 2. DSM performance is highly dependent on access patterns. For broadcast-style access patterns, contention on shared memory can drastically degrade performance. 3. Kernel configuration parameters (e.g., cluster size and block size) significantly impact DSM performance, and should be carefully selected in practical applications.
8. Dynamic Programming Instructions
@@
NVIDIA offers DPX functionshttps://docs.nvidia.com/cuda/cuda-c-programming-guide/index.html##dpx from CUDA 12 onward to accelerate dynamic programming code, enhancing programming ease. On the latest Hopper architecture, these functions are hardware-accelerated. In this section, we conduct both instruction-level and application-level benchmarks to comprehensively evaluate the potential and real-world performance of DPX.
8.1. Instruction-level
@@
Fig. 12 show the throughput of the DPX functions on three tested GPUs. The latency data is provided in the supplementary materials, which are available online. Since the DPX of RTX4090 and A100 are software emulation, their performance is almost the same. What can be observed is that for relu instructions, the performance of H800 is significantly better than the other two. For 16-bit operations, H800 also has significant acceleration, up to 13 times.
@@
Figure 12. Throughput of DPX functions on different devices
8.2. Application-level
This case study explores the application of DPX instructions to the Smith-Waterman algorithm(Smith et al., 1981), a classic dynamic programming algorithm used in bioinformatics. The Smith-Waterman algorithm finds optimal local alignments between two sequences and has wide-ranging applications in sequence alignment, genome analysis, and protein structure prediction. It employs dynamic programming by filling a two-dimensional matrix to compute the optimal local alignment score. Each matrix element represents the optimal local alignment score at corresponding positions within the two sequences. The matrix is filled recursively using a specific formula, ultimately yielding the optimal alignment.
@@ 8.2. Application-level
Our testing methodology is based on (Schmidt et al., 2024). We use the same code, data, and testing parameter configuration as described in (Schmidt et al., 2024). We use 20 protein sequences, with lengths ranging from 144 to 5478, as queries against a simulated databases. The database contains one million sequences, with fixed lengths 1024. Our results represent the average performance across all twenty queries. The performance metric for our experiments is GCUPS (Giga Cell Updates Per Second), representing the number of dynamic programming matrix cell updates performed per second.
Figure 13. Performance of the Smith-Waterman algorithm across various data types
@@
Insight 1. More complex operations (e.g., those with more operands and ReLU operations) benefit more significantly from DPX instructions. 2. The acceleration effect of 16-bit DPX instructions is more obvious, while some 32-bit operations have no acceleration effect.
9. Conclusion
This study presents a comprehensive evaluation of NVIDIA’s Hopper GPU architecture through multi-level benchmarking, revealing significant performance improvements and optimization opportunities. Our results demonstrate that Hopper’s memory subsystem achieves 2.6× higher L2 cache throughput compared to Ampere, while its innovative Distributed Shared Memory (DSM) reduces inter-SM communication latency by 6.1×. These enhancements are critical for memory-bound workloads in AI and high-performance computing.
The fourth-generation tensor cores show substantial gains when properly utilized. While backward-compatible mma instructions achieve only 62.9% of peak performance on average, the dedicated wgmma instructions fully unlock Hopper’s computational potential. Furthermore, the FP8 data type delivers 2× higher throughput than FP16 for large-scale operations, though its benefits diminish for smaller workloads due to conversion overhead.
Specialized hardware units like the TMA and DPX instructions further enhance performance. TMA enables efficient asynchronous execution, accelerating matrix multiplication by 50%, while DPX instructions improve dynamic programming workloads by up to 13× in microbenchmarks. Real-world applications, such as the Smith-Waterman algorithm, benefit significantly, achieving a 4.75× speedup with 16-bit integers.
These findings provide actionable insights for developers. To maximize Hopper’s performance, we recommend prioritizing wgmma over mma instructions, adopting FP8 for large-scale computations, leveraging DSM for efficient inter-SM communication, and utilizing DPX for dynamic programming tasks. Future work should explore optimal DSM configurations, precision-aware DPX optimizations, and the integration of TMA with other asynchronous execution features. Our research improves understanding of the latest architecture’s characteristics and performance, aiding in optimized algorithm design and application performance.
Acknowledgements.